If 6 men can do a piece of work in 4 days, how many men are needed to do the work in 21 days?
Answer by ankor@dixie-net.com(22740)

If 6 men can do a piece of work in 4 days, how many men are needed to do the work in 21 days?
:
Since they are doing an equal job
21x mandays = 4*6 mandays
:
21x = 4*6
21x = 24
x = 24/21
x = 1.143 men
:
1 man if he skips a few coffee breaks